### v1

# Flexible Music-Conditioned Dance Generation with Style Description Prompts

Abstract
Dance plays an important role as an artistic form and expression in human culture, yet the creation of dance remains a challenging task. Most dance generation methods primarily rely solely on music, seldom taking into consideration intrinsic attributes such as music style or genre. In this work, we introduce Flexible Dance Generation with Style Description Prompts (DGSDP), a diffusion-based framework suitable for diversified tasks of dance generation by fully leveraging the semantics of music style. The core component of this framework is Music-Conditioned Style-Aware Diffusion (MCSAD), which comprises a Transformer-based network and a music Style Modulation module. The MCSAD seemly integrates music conditions and style description prompts into the dance generation framework, ensuring that generated dances are consistent with the music content and style. To facilitate flexible dance generation and accommodate different tasks, a spatial-temporal masking strategy is effectively applied in the backward diffusion process. The proposed framework successfully generates realistic dance sequences that are accurately aligned with music for a variety of tasks such as long-term generation, dance in-betweening, dance inpainting, and etc. We hope that this work has the potential to inspire dance generation and creation, with promising applications in entertainment, art, and education.
Index Terms:
Pedestrian detection, hard example mining.
I Introduction
Music has always been a pivotal force in driving dance movement, whether through the rhythmic beats of traditional folk dances or the innovative choreography of contemporary performances. With the continuous development of artificial intelligence technology, there has been a growing interest in exploring the intersection between music and dance. Automated dance generation has garnered immense interest, offering fresh perspectives on the intricate relationship between music and dance, while simultaneously inspiring innovative dance creation and synthesis.
Music-conditioned dance generation aims to automatically produce dance movements that are in harmony with musical inputs. It is closely intertwined with human motion synthesis [1], which involves creating realistic and natural human motions. The main distinction lies in the fact that dance generation, which necessitates a profound analysis of music’s characteristics, is a cross-modal generation task. This task is also a typical problem in generating human motions based on conditional signals [2], which is extremely challenging due to the implicit relationship between human motion and conditional signals.
Figure 1: The illustration of dance generation with style description prompts.
Recently, advanced generative models such as Generative Adversarial Networks (GANs) [3], Variational Autoencoders (VAEs) [4], Cross-modal Transformer [5] and Diffusion [6] are employed to address music-conditioned dance generation. Although these models are capable of successfully generating high-quality and diverse dances from music signals, they overlook the significance of inherent music attributes, such as style or genre, in the dance generation process. Indeed, choreographers consider the music style to create a dance that is aesthetically pleasing to human beings. Music styles set the tone for the creative direction and style of dance generation. Different genres of music, such as classical and pop, offer varied dance forms. Therefore, careful incorporating music style is crucial in music-conditioned dance generation.
Another important aspect that cannot be neglected in dance generation is the flexibility. Dance, as a dynamic art form, often requires meticulous adjustments and refinements throughout its creation and performance phases. The ability to flexibly edit dances can not only enhance artistic expression but also enable dancers to respond effectively to user feedback. A practical dance generation approach ought to possess sufficient flexibility to cater to a diverse range of tasks.
To address above issues, we introduce a novel Music-Conditioned Dance Generation with Style Description Prompts (DGSDP), as illustrated in Figure 2. The DGSDP mainly comprises Music-Conditioned Style-Aware Diffusion (MCSAD) and Style Modulation with Description Prompts. The MCSAD is a transformer-based diffusion framework that seemly integrates music conditions and style description prompts into the dance generation. To fully utilize the style information for dance generation, we introduce a Style Modulation module and explore three types of style embeddings: one-hot encoding, genre embedding, and style description prompts. The proposed diffusion model is jointly trained with classifier-free guidance. To accommodate different dance generation tasks, a spatial-temporal masking is integrated in the backward diffusion process. To better align practical applications and facilitate future research, we design a range of experimental settings and establish benchmarks of different approaches under these settings.
The main contributions are as follows: (1) We propose flexible music-conditioned Dance Generation with Style Description Prompts (DGSDP) which deliberately incorporates style description prompts to enhance dance generation with styles. (2) We introduce Music-Conditioned Style-Aware Diffusion (MCSAD) which primarily comprises a Transformer-based dance generation network and a novel style modulation module. (3) To accommodate diverse dance generation in practical scenarios, we design a flexible dance generation algorithm by leveraging a spatial-temporal masking strategy. (4) The proposed method exhibits state-of-the-art performance across various dance generation tasks, encompassing long-term dance generation, dance in-betweening and dance inpainting.
II Related Work
II-A Stochastic Human Motion Prediction
Human Motion Prediction (HMP) can be categorized into deterministic prediction and stochastic prediction. Stochastic HMP has become popular with the emergence of generative models such as Variational Auto-Encoders (VAE) and Diffusion. Early work mostly relies on autoregressive models, which have the disadvantage of error accumulation. Yan $et$ $al.$ [7] propose a network that can generate the entire sequence directly by transforming from a series of latent vectors sampled from a Gaussian process (GP). Considering the inherent stochasticity in future motions, Kundu $et$ $al.$ [8] present a new probabilistic generative technique by introducing a random extrinsic factor. Aliakbarian $et$ $al.$ [9] propose a model with the ability to produce multiple probable future pose sequences by forcing the model to consider the noise. Mao $et$ $al.$ [10] develop a unified deep generative network and propose a pose prior and a joint angle constraint for producing smooth pose sequences. Instead of defining a combined loss as in the proposed work, Ma $et$ $al.$ [11] achieve the target of balancing the diversity sampling and the accuracy sampling by presenting a unified multi-objective conditional variational autoencoder. Given action labels and motion history, Mao $et$ $al.$ [12] design a VAE-based model which bridges the gap between stochastic human motion prediction and motion synthesis. Diffusion models have been applied for HMP. Chen $et$ $al.$ [13] present a diffusion-based framework with masked completion. Wei $et$ $al.$ [14] propose a diffusion probabilistic model to achieve diversity by incorporating a new noise at each diffusion timestep.
II-B Music-Conditioned Dance Generation
Generating dances that match the music temporally and aesthetically is challenging, and various models such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), Transformers, and Diffusion are employed to tackle this task. Li $et$ $al.$ [5] present a model involving a deep cross-modal transformer block with full-attention. Valle-Pérez $et$ $al.$ [15] introduce an innovative probabilistic autoregressive framework, taking into account prior poses and musical context. Au $et$ $al.$ [16] propose a motion generation framework, which has the capability to choreograph high-quality dance movements over a Dynamic Graph. Huang $et$ $al.$ [17] construct a transformer-based framework to generate dance with a specific genre. Kim $et$ $al.$ [3] propose a transformer-based conditional GAN framework by sampling from each latent representation of different dance genres. To generate plausible long-term dance sequences, Li $et$ $al.$ [18] introduce a two-stage process: a key pose generation and an in-between parametric motion curve prediction. Li $et$ $al.$ [19] propose a architecture with two components: a choreographic memory for encoding and quantizing dance poses and an actor-critic GPT for composing units to dance. Tseng $et$ $al.$ [6] utilize a diffusion model which enables powerful editing capabilities. Alexanderson $et$ $al.$ [20] use diffusion models for audio-driven human motion generation, using Conformers. Recently, Gong $et$ $al.$ [21] incorporate both text and music for dance generation by using a cross-modal transformer to encode multi-modal features. Unlike the above approaches, we concentrate on flexible music-conditioned dance generation and enhance generated results by integrating style information within a diffusion framework.
II-C Text-Conditioned Motion Synthesis
Text is often employed as instructions for generating motions, and similarly, it can be used as an additional condition to guide the generation of dance. Ahuja $et$ $al.$ [22] introduce a model that learns a joint embedding space of pose and natural language. Instead of long natural language, Guo $et$ $al.$ [23] propose a novel VAE framework to iteratively generate human motion sequences given a prescribed action type. Then, action is also used by Petrovich  [24], who address the problem of generating human motion sequences conditional on actions. Models in previous work all focus on an entire body. Ghosh $etal.$ [25] explore a joint mapping between 3D pose sequences and textual descriptions by presenting a hierarchical two-stream sequential model. Guo $et$ $al.$ [26] use a two-stage approach: text2length sampling and text2motion generation to generate plausible motions. Guo $et$ $al.$ [27] propose a motion token representation, which can be used in text2motion and motion2text tasks. Instead of synthesizing deterministic motions, Petrovich $et$ $al.$ [28] use variational autoencoder and a text encoder to produce diverse 3D human motions. Inspired by the powerful generative capabilities of the Diffusion model, Guy $et$ $al.$ [29] introduce a classifier-free diffusion-based generative model, which achieves impressive results in text-to-motion tasks. Zhang $et$ $al.$ [30] investigate a simple conditional generative framework based on VQ-VAE and GPT from textual descriptions. Zhang $et$ $al.$ [31] propose a diffusion model-based framework with multi-level manipulation, which can generate arbitrarily long motions and respond to fine-grained instructions on joints.
III Preliminaries
III-A Human Motion Synthesis
For human motion synthesis, auxiliary losses are frequently employed to enhance physical realism when realistic simulations are not available [32, 24]. To encourage natural and coherent motion prediction and prevent artifacts, Tevet $et$ $al.$ [29] incorporates three geometric losses: the basic joint position loss ${\mathcal{L}}_{j}$, the velocity loss ${\mathcal{L}}_{v}$ and the foot contact consistency loss ${\mathcal{L}}_{f}$. The definitions of these losses are as follows:
${\mathcal{L}}_{j}$ $=\frac{1}{N}\sum\limits_{i=1}^{N}{||\mathrm{FK}(x^{(i)})-\mathrm{FK}({\hat{x}}^{(i)})||}^{2}_{2},$ (1)
${\mathcal{L}}_{v}$ $=\frac{1}{N-1}\sum\limits_{i=1}^{N-1}{||(x^{(i+1)}-x^{(i)})-({\hat{x}}^{(i+1)}-\hat{x}^{(i)})||}^{2}_{2},$ (2)
${\mathcal{L}}_{f}$ $=\frac{1}{N-1}\sum\limits_{i=1}^{N-1}{||(\mathrm{FK}(\hat{x}^{(i+1)})-\mathrm{FK}({\hat{x}}^{(i)}))\cdot\hat{f}^{(i)}||}^{2}_{2},$ (3)
where $\mathrm{FK}(\cdot)$ represents the forward kinematic function, which transforms joint angles into joint positions, $\hat{f}^{(i)}$ the predefined binary foot contact mask, and the superscript $(i)$ indicates the frame index. Incorporating the contact consistency loss has been shown to considerably enhance the authenticity of generated motions [29].
III-B Diffusion-Based Dance Generation
Following the definition of DDPM [33], diffusion can be described as a Markov noise process, where the latent variable ${\{\sigma_{t}\}}^{T}_{t=0}$ follows a forward noise process $q(\sigma_{t}|x)$ and $x\sim p(x)$ is taken from the data distribution. The forward noise process is defined as:
$q(\sigma_{t}|x)=\sqrt{\overline{\alpha}_{t}}x+(1-\overline{\alpha}_{t})\epsilon,$ (4)
where $\epsilon\sim\mathcal{N}(0,\boldsymbol{I})$ and $\overline{\alpha}_{t}\in(0,1)$ are constants that are monotonically decreasing.
Given the music condition $c$, the model learns to estimate the ground-truth human motion $\hat{x}_{\theta}(\sigma_{t},t,c)\approx x$ at all time moments using the model parameter $\theta$ through a backward diffusion process. We optimize $\theta$ using the simple objective loss [33]:
${\mathcal{L}_{d}={\mathbb{E}}_{x,t}[{||x-{\hat{x}}_{\theta}(\sigma_{t},t,c)||}^{2}_{2}]}.$ (5)
During each denoising time step $t$, the network $\hat{x}(\cdot)$ predicts the denoised samples and adds noise $\hat{\sigma}_{t-1}\sim q(\hat{x}_{\theta}(\hat{\sigma}_{t},t,c),t-1)$, the process is repeated from $T$ to 00 and terminates when $t=0$. We employs a classifier-free guidance [34] for training, a technique commonly utilized in diffusion-based models [35, 36, 37, 29, 31]. Following the approach [34], we incorporate classifier-free guidance by introducing a low probability (e.g. 20%) of randomly replacing $c=\emptyset$ during training. The guided inference is formulated as a weighted combination of unconditional and conditional generated samples:
$\widetilde{x}(\hat{\sigma}_{t},t,c)=w\cdot\hat{x}(\hat{\sigma}_{t},t,c)+(1-w)\cdot\hat{x}(\hat{\sigma}_{t},t,\emptyset),$ (6)
where $w$ is the guidance weight with a positive value. The influence of condition $c$ can be amplified by setting $w>1$.
The overall training loss consists of a weighted combination of the primary simple loss and the supplementary auxiliary loss:
${\mathcal{L}}={\mathcal{L}}_{d}+\lambda_{j}{\mathcal{L}}_{j}+\lambda_{v}{\mathcal{L}}_{v}+\lambda_{f}{\mathcal{L}}_{f},$ (7)
where $\lambda_{j}$, $\lambda_{v}$ and $\lambda_{f}$ are weighted coefficients.
Figure 2: The proposed Dance Generation with Style Description Prompts (DGSDP) pipeline. The model is fed with a noisy motion sequence $X_{T}$ of length $N$ in a noising step $T$, along with conditioning music $c$, style description prompts $s$ and $T$ itself. The music embedding serves as input for the cross-attention module. During the inference process, DGSDP samples noise $X_{T}$ given the conditions $c$ and $s$ to predict the clean sample $\hat{X}$. Subsequently, it diffuses this sample back to $X_{T-1}$, and repeats this iterative process until $t=0$ is reached.
IV Method
We introduce a novel Dance Generation with Style Description Prompts (DGSDP), which is illustrated in Figure 2. The DGSDP mainly includes Style Modulation with Description Prompts and Style Modulation and Music-Conditioned Style-Aware Diffusion. Enhanced by a spatial-temporal masking strategy, it offers flexibility in addressing diverse dance generation tasks and possesses the capability to generate long-term dances of arbitrary lengths.
IV-A Style Description Prompts
One-Hot Encoding: One-hot encoding is commonly used to handle categorical variables, particularly when the number of categories is limited. As the music style can be divided into categories, we use one-hot encoding to represent the style information.
Genre Embedding: To obtain a more semantic representation, we utilize CLIP [38] to extract features for the words associated with the music genre, and refer to it as Style Type Embedding. This embedding captures deep contextual relationships within the textual genre, resulting in more accurate feature representations.
Style Description Prompts: Large-scale language models pre-trained on massive text [39] demonstrate impressive capabilities for text generation tasks [40]. To get a more detailed description about characteristics of the music style, we use GPT-3 [41] to generate style description prompts for dance. We choose the following function to acquire the style description prompts: ”Please generate a detailed description of the dance [$g$], including the characteristics of the dance in terms of body movement.”, where $g$ is the dance genre.
IV-B Music-Conditioned Style-Aware Diffusion
We extend the Human Motion Diffusion [33] by incorporating style modulation with description prompts, resulting in Music-Conditioned Style-Aware Diffusion (MCSAD). In addition to the music condition $c$, our model also includes the style description prompts $s$. The training objective is:
${\mathcal{L}}_{d}={\mathbb{E}}_{x,t}[{||x-{\hat{x}}_{\theta}(\sigma_{t},t,c,s)||}^{2}_{2}],$ (8)
where $\hat{x}(\cdot)$ denotes the transformer-based decoder block, which primarily comprises self-attention, feature-wise linear modulation (FiLM) [42], cross-attention, and Style Modulation.
The Style Modulation layer aims to increase the influence of style features on the generated dances. It draws inspiration from the adaptive instance normalization technique utilized in StyleGAN [43, 44]. This module $\mathrm{SM}(\cdot)$ allows for the incorporation of additional textual information into existing transformer models. Formally,
$\mathrm{SM}(z,s)=\frac{z}{{||z||}_{2}\cdot r}\cdot\mathrm{FC}(s),z\in{\mathbb{R}}^{T\times d_{z}},s\in{\mathbb{R}}^{d_{s}},$ (9)
where $z$ refers to the input, $s$ represents the music style description prompts, $r$ is a scaling factor, $\mathrm{FC}(\cdot)$ denotes the fully connected layer, $d_{s}$ is the embedding dimension and $d_{z}$ is the hidden dimension.
With the style description prompts, the guided inference is:
$\widetilde{x}(\hat{\sigma}_{t},t,c,s)=w\cdot\hat{x}(\hat{\sigma}_{t},t,c,s)+(1-w)\cdot\hat{x}(\hat{\sigma}_{t},t,\emptyset,s).$ (10)
The detailed training algorithm is shown in Algorithm 1.
Require : The initialized dance generation network $\hat{x}_{\theta}$, the noising steps $T$, the maximum iterations $I_{max}$.
Input : The known dance sequences $x$, the music conditioning $c$ and style description prompts $s$.
Output : The dance generation network $\hat{x}_{\theta}$.
for $I=0,1,...I_{max}$ do
       Sample the timestep $t=Uniform({1,2,...,T})$.
      Randomly sample a noise $\epsilon\sim\mathcal{N}(0,\boldsymbol{I})$.
      Add noise on $x$ with $\epsilon$ and pre-defined parameter $\overline{\alpha}_{t}$: $\sigma_{t}=\sqrt{\overline{\alpha}_{t}}x+(1-\overline{\alpha}_{t})\epsilon$.
      Replace $c$ with $\emptyset$ with a low probability.
      Calculate the losses $L_{j}$, $L_{v}$, and $L_{f}$ defined in Eq. (1-3).
      Calculate the training loss $L_{d}$ defined in Eq. (8).
      Update parameters of the network: $\theta=\theta-\nabla_{\theta}({\mathcal{L}}_{d}+\lambda_{j}{\mathcal{L}}_{j}+\lambda_{v}{\mathcal{L}}_{v}+\lambda_{f}{\mathcal{L}}_{f})$.
end for
Algorithm 1 Training Algorithm of our Pipeline
Figure 3: Flexible dance generation with spatial-temporal masking. For the known sequence, we add noise on it to obtain noisy sequence at timestep $t-1$ directly. For unknown sequences, we first use the trained network ${\hat{x}}_{\theta}$ to predict the motion at the timestep 0, and then we add noise on it to obtain noisy sequence at timestep $t-1$ The mask is two-dimensional and allows for control in both the temporal dimension (frame) and spatial dimension (joint).
IV-C Flexible Generation with Masking
During inference, the proposed method generates an estimated sequence from the initial noisy sequence $\sigma_{T}\sim\mathcal{N}(0,\boldsymbol{I})$, denoises it to $\hat{\sigma}_{T-1}$, and iterates this process until $t=0$.
To flexibly edit generated dances, we design a spatial-temporal masking strategy, which is illustrated in Figure 3. We first pad the known dance sequence to match the size of the target generation. Then, for the padded known dance sequence $x_{0}$, we add noise on it to obtain noisy sequence at the timestep $t-1$,
$x^{known}_{t-1}=\sqrt{\overline{\alpha}_{t-1}}x_{0}+(1-\overline{\alpha}_{t-1})\epsilon,$ (11)
where $\epsilon\sim\mathcal{N}(0,\boldsymbol{I})$.
For an unknown sequence, the model initially predicts the target sequence from random noise, subsequently introducing noise to generate a noisy sequence at the timestep $t-1$,
$x^{unknown}_{t-1}=\sqrt{\overline{\alpha}_{t-1}}\hat{x}_{\theta}(\hat{\sigma}_{t},t,c,s)+(1-\overline{\alpha}_{t-1})\epsilon.$ (12)
Finally, $x^{known}_{t-1}$ and $x^{unknown}_{t-1}$ are combined with a spatial-temporal mask $M$,
${\hat{\sigma}}_{t-1}=M\odot x^{known}_{t-1}+(1-M)\odot x^{unknown}_{t-1},$ (13)
where the mask $M\in{\{0,1\}}^{F\times J}$ can be changed arbitrarily with 0 or 1 values in both the temporal and spatial dimensions, $F$ and $J$ represent the number of frames and joints, respectively. This process is iterated by taking ${\hat{\sigma}}_{t-1}$ as the noise for the next iteration. The spatial-temporal mask supports any combination of temporal constraints and joint constraints. This masking framework offers a robust tool for subsequent applications to generate dance sequences that precisely adhere to arbitrary constraints. Details of the algorithm are shown in Algorithm 2.
Require : The diffusion timestep $T$, music conditioning $c$, style description prompts $s$ and the trained dance generation network $\hat{x}_{\theta}$.
Input : The padded dance sequences $x_{0}$, corresponding masks $M\in{\{0,1\}}^{F\times J}$.
Output : The completed dance sequences $x$.
Randomly sample a noise $\hat{\sigma}_{T}\sim\mathcal{N}(0,\boldsymbol{I})$;
for $t=T,T-1,...1$ do
       Randomly sample a noise $\epsilon\sim\mathcal{N}(0,\boldsymbol{I})$ if $t>1$, else $\epsilon=0$.
      For the known sequences $x_{0}$, add noise on them with $\epsilon$ and pre-defined parameter $\overline{\alpha}_{t-1}$: $x^{known}_{t-1}=\sqrt{\overline{\alpha}_{t-1}}x_{0}+(1-\overline{\alpha}_{t-1})\epsilon$.
      For the unknown sequences, first use pre-trained network $\hat{x}_{\theta}$ to reconstruct $\hat{x_{0}}$ and then add noise on it with $\epsilon$ and pre-defined parameter $\overline{\alpha}_{t-1}$: $x^{unknown}_{t-1}=\sqrt{\overline{\alpha}_{t-1}}\hat{x}_{\theta}(\hat{\sigma}_{t},t,c,s)+(1-\overline{\alpha}_{t-1})\epsilon$.
      Combine $x^{known}_{t-1}$ and $x^{unknown}_{t-1}$ with a spatial-temporal mask $M$: ${\hat{\sigma}}_{t-1}=M\odot x^{known}_{t-1}+(1-M)\odot x^{unknown}_{t-1}$.
end for
Algorithm 2 Flexible Dance Generation Algorithm with Masking
The proposed framework can also be applied for long-term dance generation. Since the model generates a set of frames of a dance sequence at once, increasing the maximum sequence length results in a linear increase in computational cost. In addition, dance generation requires the music condition to match the motion sequence in length, which further extends the memory requirements. We follow the long-term generation strategy [6] to achieve temporal consistency between multiple sequences in long-term generation.
V Experiments
V-A Dataset and Implementation Details
The AIST++ dataset [5] includes 1,408 high-quality dance movements synchronized with music from 10 genres. It comprises motion sequences of varying durations, ranging from 7 seconds to 50 seconds, with an average duration of 13 seconds. We follow the original train/test splits. Consistent with EDGE [6], instances in the training set are cut to 5 seconds at 30 FPS with a stride of 0.5 seconds, and instances in the testing set are cut to 5 seconds at 30 FPS with a stride of 2.5 seconds.
The proposed model has 49.7 million (M) parameters, and was trained on one NVIDIA RTX A6000 GPU for 2 days with a batch size of 128. The implementation is based on the PyTorch. The number of iterations is 1000. The learning rate is 0.0002 and the weight decay is 0.02. The music condition transformer encoder has a depth of 2 layers, with 8 attention heads and a dimension of 512. The decoder block has a depth of 8 layers, with 8 attention heads and a dimension of 512. For long-term generation, we use linear weighted summation to sum up corresponding parts of different slices (5s). Specifically, we perform a weighted summation of the latter half (2.5s) of the previous slice and the former half (2.5s) of the next slice. To mitigate the impact of randomness of stochastic, all evaluation metrics obtained by averaging over 100 trials.
Following [19, 6], we adopt the following evaluation metrics:
Beat Alignment Score: It is used to measure the alignment between the music and the generated movement. We calculate the average time distance between each music beat and its closest dance beat as the beat alignment score. The music beats are extracted using the Librosa library [45], and the kinematic beats are calculated based on the local minima of the velocity of motion joints.
Physical Foot Contact Score: It is a quantitative acceleration-based metric used to score the physical plausibility of the generated kinematic motion. The metric derives from two simple and related observations: (1) In the horizontal plane, any centre of mass (COM) acceleration must be due to static contact between the feet and the ground. Thus, either at least one foot is stationary on the ground or there is no acceleration of COM. (2) On the vertical axis, any positive COM acceleration must be due to static foot contact.
Frechet Inception Distance(FID): FID is the distance between the generated dance movement distribution and the real movement distribution. We calculate the FIDs between the generated dances and the AIST++ dataset for all motion sequences in the kinetic feature space [46] (denoted as $FID_{k}$) and geometric feature space [5] (denoted as $FID_{g}$).
Diversity: This metric assesses the diversity of dance motions by computing the average Euclidean distance between different generated dance motions. Similarly, we compute the Diversity between all motion sequences of the generated dances in the kinetic feature space [46] (denoted as $Div_{k}$) and the geometric feature space [5] (denoted as $Div_{g}$).
TABLE I: Comparision of dance generation on the AIST++ dataset. For simplicity, $\uparrow$ means higher is better, $\downarrow$ means lower is better and $\rightarrow$ means closer to the ground truth is better. During inference, akin to EDGE [6], the guidance weight $w$ is set to 2 by default, and the results obtained with $w=1$ are also reported to explore the influence of amplified conditions.
Method Guidance weight $w$ = 2 Guidance weight $w$ = 1 $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ EDGE [6] 0.26 1.56 35.35 18.92 5.32 4.90 0.25 1.19 45.41 19.42 4.82 5.24 DGSDP (ours) 0.31 1.51 34.32 19.04 6.29 5.81 0.28 1.65 37.21 18.19 7.25 6.42 Ground Truth 0.35 1.33 9.43 7.33 0.35 1.33 9.43 7.33
TABLE II: Results of long-term dance generation on the AIST++ dataset.
Method 7.5 seconds 10 seconds $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$ EDGE [6] 0.26 1.11 59.43 25.60 2.97 3.52 0.25 0.88 68.63 32.24 2.55 3.14 DGSDP (ours) 0.30 1.38 55.65 20.74 4.10 5.74 0.30 2.13 56.15 29.98 5.21 6.45 Ground Truth 0.38 1.04 9.34 7.47 0.49 1.70 9.34 7.47
V-B Experimental Settings
We evaluate the performance of flexible dance generation in the following scenarios.
Dance Generation: These setting generates dance movements of 5-second length given the conditioned music. This is the standard setting for music-conditioned dance generation.
Long-Term Dance Generation: This setting aims to generate 7.5-second and 10-second dance movements. For fair comparison, the model is trained on 5-second clips.
Seed Motion: Given the real motions of the first and second frames, the model generates dance movements of the following frames.
In-betweening: Given the real motions of the first and last frames, the model completes the motions of the in-between frames.
Inpainting: Given a random mask, the model completes the remaining frames based on the mask. We control the model to randomly give 70% of the true sequences.
Upper-body Generation: Since the mask possesses both spatial and temporal dimensions, apart from controlling it in the temporal dimension, we can also manipulate it in the spatial dimension. Upper-body generation aims to generate motions of the upper-body given lower-body motions.
Lower-body Generation: Given upper-body dance movements of all sequences, the model automatically generates lower-body dance movements.
V-C Results of Dance Generation
We choose EDGE [6] as our baseline. For different approaches, we generate 20 pieces of dances with 5-second length, covering 10 different genres of AIST++. Quantitative results are summarized in Table I. Our method performs favorably against EDGE [6] on most metrics. Specifically, there is a significant improvement in the Beat Alignment Score metric, which is main metric of dance generation, suggesting that the dances generated of our method are more in tune with the musical melody. Moreover, our method achieves the lowest $FID_{k}$, which is defined on motion velocities and energies, reflecting the physical characteristics of dance. The superiority of $FID_{k}$ demonstrates that the proposed DGSDP can generated more realistic dance movements. In addition, our method achieves the closest approximation to the ground-truth values of $Div_{k}$ and $Div_{g}$, revealing the capability to generate more diverse dance movements of our method rather than converging to a few fixed future motions.
V-D Results of Long-Term Dance Generation
Since both the training and test sets are composed of five-second slices, in order to validate the ability of our model to generate long-term dance sequences, we generate the corresponding dances for music that are 7.5 seconds and 10 seconds. We select the pieces of dances that are longer than 7.5 seconds or 10 seconds for long-term dance generation. Experimental result of long-term dance generation on AIST++ is shown in Table II.
Our method performs significantly better than EDGE [6] on most metrics. Specifically, our model achieves a higher Beat Alignment Score, maintaining good performance on the correlation between dance movements and music. What’s more, our model performs better on EDGE [6] on metrics $FID_{k}$,$FID_{g}$,$Div_{k}$ and $Div_{g}$. This reveals that DGSDP can also generate dance movements with better quality and better diversity in long-term dance generation tasks. It is worth noting that the performance of our model on $Div_{k}$ and $Div_{g}$ do not decrease significantly over time like EDGE [6]. This suggests that, despite for long music, the dance movements generated using our method can ensure diversity of movements, avoiding repetitive movements.
V-E Results of Flexible Dance Generation
With the spatial-temporal masking strategy, our approach is applicable to generate dance movements according to various scenarios. We quantitatively compare results of ours with those of EDGE [6] for flexible dance generation in Table III.
In the seed motion scenario and upper-body generation scenario, our method performs favorably against EDGE [6] on all metrics. In the other three scenarios, our method performs better than EDGE [6] on most metrics. This suggests that our approach has a strong capacity for flexible generation from both the temporal and spatial perspectives. Furthermore, it is clearly found that the performance of our model on Beat Alignment Score is close to the ground truth or even exceeds it, revealing the potential for strong self-improvement properties of our model in harmonising music and movement. Overall, our approach is suitable for a wide range of flexible scenarios, and with the help of known information, our model may perform better.
TABLE III: Results of flexible dance generation.
Scenarios Method $BeatAlign\uparrow$ $PFC\downarrow$ $FID_{k}\downarrow$ $FID_{g}\downarrow$ $Div_{k}\rightarrow$ $Div_{g}\rightarrow$
Seed motion EDGE [6] 0.27 1.19 32.62 21.68 5.81 5.15
DGSDP (ours) 0.33 1.18 27.05 17.36 7.34 5.79
In-betweening EDGE [6] 0.28 1.17 37.33 15.16 4.58 4.86
DGSDP (ours) 0.34 1.24 30.20 14.75 5.80 6.16
Inpainting EDGE [6] 0.36 2.15 21.57 17.50 7.47 8.36
DGSDP (ours) 0.37 1.71 23.80 18.78 7.03 8.34
Upper-body generation EDGE [6] 0.29 1.67 32.90 14.82 5.05 5.45
DGSDP (ours) 0.33 1.40 32.86 14.36 5.39 6.11
Lower-body generation EDGE [6] 0.33 0.92 39.81 15.44 5.28 6.86
DGSDP (ours) 0.36 0.99 42.26 14.73 5.39 7.37
Ground Truth 0.35 1.33 9.43 7.33
[…]
Figure 4: Visualization of generated dances for the same piece of music. We list four music genres, each associates with five different dance movements. Generated dance movements are in blue and real movements are in grey.
V-F Discussions
We conduct detailed ablation studies, analyze the impacts of audio representations and various style embeddings, and discuss the limitations of evaluation metrics.
Audio Representation: We evaluate three independent feature extraction strategies: Jukebox [47], Encodec [48] and Librosa [49].
Jukebox [47], a pre-trained generative model designed for music generation, showcasing notable performance in tasks specific to music prediction as evidenced in prior studies [6]. Musicians and composers can use Jukebox to create original musical compositions, exploring and innovating between different genres and styles. Inspired by these advances, here, we extract Jukebox music features as music conditions.
Encodec [48], an advanced real-time, high-fidelity, audio codec. Encodec is a pre-trained model, which has been used as music conditional feature extractor in dance generation tasks as well. It achieves state-of-the-art quality levels in both voice and music compression at multiple audio compression ratios and sample rates.
Librosa [49] is a library for analyzing audio signals. It provides a rich set of functions including loading audio files, extracting features, computing rhythmic features and performing audio signal processing. Here, we use librosa to extract the following musical features: envelope (1-dim), MFCC (20-dim), chroma (12-dim), one-hot peaks (1-dim) and one-hot beats (1-dim).
[…]
It is demonstrated through ablation experiments that our method achieves the best Beat Alignment Score and maintains superior performance in other metrics.
Visualizations: For the same music, our approach can generate a variety of sensible dance movements. Some of the dance visualization results are shown in Figure 4. Here, we list generated dances for four types of music: Breaking, Krump, Locking and Waacking. For each type, we list five dance movements. As can be seen from the figure, our approach can generate diverse and plausible dance movements.
Limitations of Evaluation Metrics: Developing automated metrics for the quality of the dance generated is a challenging endeavour due to the complex, subjective and even culturally specific context of dance practice. FID-based metrics are suitable for data-rich domains such as image generation, but the AIST++ test set is too small to cover the train distribution. Moreover, because of the limited availability of data, both $FID_{k}$ and $FID_{g}$ rely on heuristic feature extractors that merely calculate surface-level characteristics of the data. We argue that the idea of assessing the difference between two distributions of features of dance movements is not necessarily fundamentally flawed, but that more representative features may lead to reliable automated quality assessments.
The beat alignment score calculates the average time distance between each music beat and its closest dance beat. However, dancing is not solely about syncing local speed minima in joints with beats. Instead, musical beats serve as a flexible reference for timing, rhythm, and the smooth transition between dance movements and steps. While this metric has driven progress on this issue in the past, more dance-specific metrics are needed to realize breakthroughs in this area.
VI Conclusion
In this work, we study music-conditioned dance generation and introduce a transformer-based diffusion framework called Dance Generation with Style Description Prompts (DGSDP) which can generate arbitrary dance sequences. The DGSDP mainly comprises Music-Conditioned Style-Aware Diffusion (MCSAD) and Style Modulation with Description Prompts. Our approach obtains state-of-the-art performance on a variety of tasks such as dance generation, long-term dance generation, dance in-betweening, dance inpainting. Ablated experiments demonstrate the effectiveness of the modulation module and the style description prompts. We also conduct a thorough investigation into various encodings of both textual music style and the audio music. To obtain dances that align better with the music, we suggest using Jukebox to extract audio representations and applying style description prompts augmented by large language models. The proposed framework is suitable for flexible dance generation, enabling the user to freely edit or generate the desired dance sequence. We hope that this work will inspire further research in the fields of automatic dance generation and interactive dance generation.